OpenStudy (anonymous):

use substitution to solve the following system of equations. d+e-f=11 e=f+d+5 f=2e-12 ps. if solve you get a medal

OpenStudy (anonymous):

solve one of the equations for one of its variables, it doesn't matter which one you chose because there is no incorrect choice... f = 2e - 12 now substitute the value from the first variable you solved for into any of the other equations and solve for the next variable e = f + d + 5 e = (2e - 12) + d + 5 e = 2e + d - 7 -e = d - 7 e = -d + 7 now substitute the two variables you have into the last equation... d + e - f = 11 d + (-d + 7) - (2e - 12) = 11 d - d + 7 - 2e + 12 = 11 19 - 2e = 11 19 - 11 = 2e 8 = 2e 4 = e now sub 4 in for e... e = -d + 7 4 = - d + 7 -3 = - d 3 = d now sub in your known variables into original equation to solve for last variable... d + e - f = 11 3 + 4 - f = 11 7 - f = 11 -f = 4 f = - 4 now check... f = 2e - 12 -4 = 2(4) - 12 -4 = 8 - 12 -4 = -4 (correct) answer : e = 4, f = -4, d = 3 If you have any questions at all about what I did, please ask. you can also look here.... http://www.mrferrante.com/algebra2/final_review/chapter3/3.6.1.html